OpenStudy (anonymous):

men's white handkerchief costs 2.29 for 3. the cost per dozen is

OpenStudy (anonymous):

good way to look at these; especially when you have a lot of factors to consider: \[\frac{ $2.29 }{ 3 handkerchiefs } \times \frac{ 12 handkerchiefs }{ 1 dozen }\] the word/variable handkerchief cancels out leaving \[\frac{ $2.29 }{ 3 } \times \frac{ 12 }{ 1 dozen } \]

OpenStudy (anonymous):

= 2.29 x 4

OpenStudy (anonymous):

thanx